OpenStudy (anonymous):

here is my problem, P=3000, A=$3180, t=1 year, how do I find the rate?

OpenStudy (anonymous):

the rate of interest?

OpenStudy (studentchaitrali):

Is it compound interest or simple interest?

OpenStudy (studentchaitrali):

Compound interest: A=P((100+R)/100)^T Simple interest: SI=PRT/100 SI=A-P

OpenStudy (wolf1728):

You need to solve the equation for rate. I have attached a formula solved for the rate.

OpenStudy (studentchaitrali):

Its 6% per annum

OpenStudy (wolf1728):

The equation is complex and is split into 2 parts. exp= [log(total/princ)/years] exp = log(3180/3000)/1 exp = log(1.06) exp= 0.0253058653 rate = (10^exp -1) rate = 10^0.0253058653 -1 rate = .06 So, it must be simple interest. If it were compounded interest you'd definitely need to use the formula.